# Leveraging Content and Acoustic Representations for Efficient Speech Emotion Recognition

### v1

@@ Abstract
Speech emotion recognition (SER), the task of identifying the expression of emotion from spoken content, is challenging due to the difficulty in extracting representations that capture emotional attributes from speech. The scarcity of large labeled datasets further complicates the challenge where large models are prone to over-fitting. In this paper, we propose CARE (Content and Acoustic Representations of Emotions), where we design a dual encoding scheme which emphasizes semantic and acoustic factors of speech. While the semantic encoder is trained with the distillation of utterance-level text representation model, the acoustic encoder is trained to predict low-level frame-wise features of the speech signal. The proposed dual encoding scheme is a base-sized model trained only on unsupervised raw speech. With a simple light-weight classification model trained on the downstream task, we show that the CARE embeddings provide effective emotion recognition on a variety of tasks. We compare the proposal with several other self-supervised models as well as recent large-language model based approaches. In these evaluations, the proposed CARE model is shown to be the best performing model based on average performance across $8$ diverse datasets. We also conduct several ablation studies to analyze the importance of various design choices.
@@ I Introduction
Speech Emotion Recognition (SER) involves identifying the emotional state of the speaker based on acoustic characteristics. The ability to recognize emotions from speech has profound applications across various domains, such as human-computer interaction [1], social media analytics [2], customer call centres [3] and mental health monitoring systems [4]. Despite significant advancements, SER remains a challenging problem due to the complex nature of human emotions and the difficulties in succinctly capturing the human emotions from limited supervised datasets.
Traditionally, SER systems used different acoustic properties of speech signals to analyze emotion content. The importance of the pitch contour towards this task was first highlighted by Lieberman et al. [5]. Other acoustic characteristics such as energy, intensity and speaking rate were subsequently established as markers of emotion [6]. These features, proposed as part of the Interspeech para-linguistic challenges, were high dimensional  [7, 8]. The dimensionality issue was partially addressed by Eyben et al. [9], where a minimalist set of features were proposed for SER. However, the hand crafted features suffer from lack of robustness to novel emotional datasets and tasks.
In the last decade, the research efforts in SER progressed from using hand crafted features to deep learning techniques. The network architectures include convolutional neural networks (CNN) [10, 11], long short term memory (LSTM) [12] networks, and transformer models [13]. More recently, self-supervised learning (SSL) has emerged as a viable alternative to the classical approach of training individual models for different tasks. These SSL models, trained on large amounts of unlabeled data, have been shown to generalize well to a variety of downstream tasks. The popular examples include wav2vec2.0 [14], HuBERT [15] and WavLM [16], which have shown promising results on emotion recognition [17, 18].
With the advent of large language models (LLMs), recent efforts in speech representation learning attempt to encode speech using speech-to-text encoders like Whisper [19] followed by large language modeling with adapters. Examples of such efforts include SALMONN [20] and wavLLM [21]. Further, the alignment of speech and text at embedding level was attempted by SONAR [22]. However, these efforts require a large model with billions of parameters to perform tasks like emotion recognition.
In this paper, we propose the design of a novel parameter-efficient modeling framework of representing speech in a dual encoding fashion. The proposed model is termed as Content and Acoustic Representations of Emotions, where one of the encoder arms, termed semantic encoder, aligns the speech representations with the recognition transcripts encoded using the RoBERTa model [23] at sentence level, while the non-semantic arm aligns the speech representations with low-level descriptors of speech, given by the PASE+ model [24]. Both the encoder arms are learned through distillation losses, with the weights initialized using the WavLM [16] SSL model. As the emotion recognition and understanding tasks entail both the semantic and non-semantic representations, we concatenate the two encoder outputs and learn a light-weight classification head to perform emotion recognition for the downstream tasks. We show the effectiveness of the CARE approach by evaluating on $8$ speech emotion datasets as well as on a depression detection dataset. The key contributions of our work are as follows:-
  • Designing a parameter efficient SER system, termed as CARE, consisting of a semantic encoder and an acoustic encoder.
  • Performing distillation based learning of the CARE model by leveraging pre-trained SSL models of speech - WavLM [16] and PASE+ [24]) for acoustic arm, Whisper [19] and text (RoBERTa [23]) models for semantic arm.
  • Proposing a convolution layer adaptation of text-based models for representation learning of speech.
  • Experimentation on $8$ benchmark speech datasets with diverse SER tasks - achieving state-of-art (SoTA) performance on some these tasks while outperforming models of much larger size.
@@ II-A Audio feature extraction for SER
Speech emotion is often conveyed through various factors encoded in speech like content, pitch, prosody, stress, loudness etc [25]. Earlier works attempted to extract wide variety of acoustic descriptors for emotion recognition [9]. Deep learning based representations have been increasingly used as low-level acoustic features. Some examples in this line of work include the SincNet architecture by Ravanelli et al. [26], problem agnostic speech representations (PASE) by Pascual et al. [27] and interpretable Gaussian filters by Agrawal et al. [28]. One such architecture, called the learnable front-end for audio classification (LEAF) [29] was used by Dutta et al. [11] for speech emotion classification. In contrast, this work designs a setting where the semantic features are also represented along with low-level acoustic features for emotion recognition.
II-B Self supervision in speech
Following the application of self-supervised models for text, a large number of SSL models emerged in speech. One of the earliest such models was proposed by Schneider at al. [30]. The technique of masking out parts of the data and subsequently predicting them, also called masked language modeling (MLM), was introduced in the domain of natural language processing [31]. A similar setting for speech was adopted by Baevski et al. [14]. An iterative training procedure, called HuBERT, proposed by Hsu et al. [15], used MLM loss and led to further improvements in various downstream tasks. The approach termed WavLM [16] explored augmenting the pre-training recipe for HuBERT with a speech separation task. Recent approaches in this direction include teacher-student learning setting [32] or disentangled learning [33, 34].
II-B1 Self supervision for SER
Siriwardhana et al. [35] proposed a multimodal model for emotion recognition wherein they processed the speech signal by a vq-wav2vec model [36] followed by a pre-trained speech-BERT [37] model. Similarly, the pre-trained wav2vec model has been used in other works [38, 39]. Pepino et al. [40] showed the effectiveness of fusing the wav2vec2.0 features along with the eGeMAPS features for emotion recognition.
We describe two recent methods in this direction - Vesper by Chen et al. [41] and emotion2vec [42]. The primary objective of Vesper was to compress a pre-trained model for emotion recognition in a teacher-student mode of training, with the teacher being initialized with the WavLM-large model. Emotion2vec [42] is a method where the data2vec [32] model is considered as the teacher which is used to train a student model. In contrast to the proposed work, the prior SSL efforts for SER largely ignore the high-level semantic content information, which is also observed to contribute significantly to emotion recognition tasks [18].
@@ II-C Speech-text aligned representations
The alignment of speech and text modalities is an increasingly active area of interest for speech representation learning. The joint pre-training of speech and text using supervised data was proposed by Bapna et al. [43]. The MAESTRO model, proposed by Chen et al. [44], explored the learning of a shared speech-text encoder and a duration model which upsampled the text embeddings to match the speech duration. In this line of work, Speech-LM, proposed by Zhang et al. [45], investigated the modeling of unpaired speech and text data. In contrast to these efforts, the CARE model attempts to jointly model textual and non-textual speech representations.
The SONAR model [22] designs a speech encoder which aligns with the encoded textual representations at the utterance level. With the rising popularity of large language models (LLMs), several recent works have developed speech encoders that are jointly learned with LLM models. Many such works use the Whisper ASR encoder [19] with LLaMA [46] LLMs. The SALMONN model by Tang et al. [20] developed speech-music encoder for broader audio understanding, while WavLLM, proposed by Hu et al. [21], used Whisper and WavLM encoders along with low-rank adaptation (LoRA) [47] of LLMs. While these LLM based efforts use speech-text aligned representations with other acoustic encoders and enable prompt based applications, the models are inherently large-sized ($8$B) with a significant burden on training and inference. In contrast, our proposed CARE system, even with its compact size of $160$M parameters, outperforms larger models in several downstream evaluations.
Figure 1: Block diagram of the proposed CARE model. An ASR system is used to generate the transcripts for the pre-training data which is passed through a pre-trained RoBERTa model to generate targets for the semantic arm. The acoustic arm of the model is trained with PASE+ features as targets. Blocks in blue indicate either frozen components or those with no learnable parameters.
@@ III-A Background
III-A1 WavLM
The MLM task in the speech domain, with architectures like wav2vec2.0 [14] and HuBERT [15], advocates a style of pre-training, wherein segments of the speech signal are masked out and the transformer based models are trained to predict the masked speech regions. The WavLM [16] uses the same pre-training architecture as HuBERT as well as the clustering based discrete tokenization of audio with the addition of a de-noising objective. The speech signals are randomly mixed with noise or with other speech signals to create noisy or overlapped speech. A stack of convolutional layers transform the speech data followed by $12$ ($24$ for WavLM-large) transformer layers to predict the discrete cluster labels corresponding to clean speech. This de-noising objective, along with the masked language modeling loss, results in improved performance on non-ASR tasks compared to HuBERT or wav2vec2.0 [16].
III-A2 RoBERTa
One of the significant contributions in creating a large scale language model was proposed by Devlin et al. [31]. This architecture, called bidirectional encoder representations from transformer (BERT), was trained with two objectives on a corpus of textual data, namely, predicting words masked out in a sentence and to predict whether two sentences semantically follow each other (referred to as next sentence prediction). Liu et. al [23] trained this architecture on a larger corpus of textual data without the next sentence prediction task. This pre-trained model is known as robust optimized BERT approach (RoBERTa). We use the pre-trained base version of this model with $12$ transformer layers as the text encoder.
@@ III-B CARE Model
The block diagram of the CARE model is shown in Fig.1. Text supervision: The annotation of text transcripts for the emotional pre-training speech datasets is highly demanding. A pre-trained ASR system offers an alternate way to generate “noisy” transcripts of speech. We use an open source ASR model, Whisper-large-v3 [19], for converting speech to text transcripts. Note that, this ASR model was not trained on emotional speech, hence the word-error-rate (WER) on emotional speech datasets is higher than those on neutral speech datasets [18, 48]. The noisy text transcripts are encoded using the RoBERTa model [23] and the mean-pooled utterance features serve as the teacher signal for the semantic arm of the CARE model.
Acoustic Supervision: The non-semantic components of speech, typically captured in low-level acoustic descriptors of speech like filter-bank energies, prosody, pitch etc are also potentially rich speech representations containing information about the underlying emotion. The PASE+ model [24] contains a series of convolutional and recurrent neural network layers which encode the speech, followed by a set of workers that regress the low-level descriptors of speech, derived from the clean input signal. We use the frame-level PASE+ features of $256$ dimensions as targets for the acoustic arm of the CARE model. These features are down-sampled by a factor of $2$ to generate targets at $50$Hz.
@@ III-B1 Model Architecture
The speech signal is initially processed by a set of convolutional layers. The stack of $6$ transformer layers acts as the common encoder arm. Both the acoustic and semantic arms are $6$-layer transformer models. They differ in the distillation targets that are used to train the respective arms. The acoustic encoder uses the PASE+ representations as targets, while the semantic encoder uses the RoBERTa representations of the ASR transcripts as targets. Further, the acoustic arm receives a frame-level loss while the semantic encoder receives an utterance level loss. The utterance level objective function for text data avoids the need for time-aligning speech and text, one of the key difficulties in training speech-text joint models [44].
For the semantic arm, we propose a novel adaptation strategy. The last $6$ transformer layers of the pre-trained RoBERTa model are used as frozen weights for the semantic arm of the CARE model. In the semantic arm, two convolutional blocks are trained, one before and one after each transformer layer. These convolutional blocks serve the purpose of adapting the speech representations to the internal representations of the RoBERTa model. The convolutional block before the transformer layer down-samples the sequence length by a factor of $3$, while the block following the transformer layer up-samples the representations by the same factor. This ensures that the frozen RoBERTa transformer layer processes sequences with lower time-resolution than the original speech. The representations from the semantic arm are average pooled to generate an utterance level representation of the speech signal.

### v2

# Leveraging Content and Acoustic Representations for Speech Emotion Recognition

@@ Abstract
Speech emotion recognition (SER), the task of identifying the expression of emotion from spoken content, is challenging due to the difficulty in extracting representations that capture emotional attributes from speech. The scarcity of labeled datasets further complicates the challenge where large models are prone to over-fitting. In this paper, we propose CARE (Content and Acoustic Representations of Emotions), where we design a dual encoding scheme which emphasizes semantic and acoustic factors of speech. While the semantic encoder is trained using distillation from utterance-level text representations, the acoustic encoder is trained to predict low-level frame-wise features of the speech signal. The proposed dual encoding scheme is a base-sized model trained only on unsupervised raw speech. With a simple light-weight classification model trained on the downstream task, we show that the CARE embeddings provide effective emotion recognition on a variety of datasets. We compare the proposal with several other self-supervised models as well as recent large-language model based approaches. In these evaluations, the proposed CARE is shown to be the best performing model based on average performance across $8$ diverse datasets. We also conduct several ablation studies to analyze the importance of various design choices.
@@ I Introduction
Speech Emotion Recognition (SER) focuses on detecting the speaker’s emotional state from the audio signal. Recognizing emotions in speech has significant applications across diverse fields, including human-computer interaction [1], social media analysis [2], customer service call centers [3], and mental health monitoring systems [4]. However, despite considerable progress, SER continues to pose challenges due to the complexity of human emotions and the inherent difficulties in effectively capturing these emotions from limited labeled datasets.
Traditionally, SER systems have relied on various acoustic properties of speech signals to interpret emotional content. Lieberman et al.[5] were among the first to emphasize the role of pitch contour in emotion analysis. Additional acoustic features, including energy, intensity, and speaking rate, were later recognized as important indicators of emotional class [6]. The features identified through the Interspeech para-linguistic challenges were rich in emotional properties while being inherently high-dimensional [7, 8]. Eyben et al. [9] introduced a minimalist feature set to address this dimensionality issue partially.
In recent years, the field of speech emotion recognition (SER) has shifted to the adoption of deep learning methods. Current network architectures in SER commonly include convolutional neural networks (CNN)[10, 11], long short-term memory (LSTM) networks[12], and transformer models [13]. While these models perform well on the specific datasets they are trained on, they often struggle to generalize across diverse datasets. Due to the difficulty of training individual models for each dataset, self-supervised learning (SSL) models have emerged as a promising solution.
Self-supervised learning (SSL) models, trained on large volumes of unlabeled data, have shown impressive generalization capabilities across a range of downstream tasks. Notable examples include wav2vec 2.0 [14], HuBERT [15], and WavLM [16]. These models are engineered to capture speech patterns in a manner similar to how models like BERT [17] capture language patterns. Although trained on neutral speech data, these models have demonstrated encouraging results in emotion recognition tasks [18, 19]. The emotion recognition performance may be additionally enhanced by training these models with tailored self-supervised objective functions that are emotion aware. Two recent examples are Vesper [20] and emotion2vec [21]. Both methods share a common strategy of utilizing a student-teacher framework for distilling knowledge from an existing model, such as WavLM (in the case of Vesper) or data2vec (in the case of emotion2vec), while training on emotional datasets. However, emotion is shaped not only by these features but also by the semantic content of the speech signal [22]. For instance, identifying emotions from text transcripts is often more effective than interpreting them from raw audio [19]. The integration of speech content during the pre-training phase of SER models remains an under-explored yet promising area of research.
In this work, we introduce a self-supervised model for speech emotion recognition (SER) called Content and Acoustic Representations of Emotions (CARE). To the best of our knowledge, our approach is the first effort to pre-train a self-supervised model that integrates both semantic and acoustic components of speech. CARE leverages a dual encoding framework for processing speech signals: a semantic encoder, which aligns speech representations with sentence-level transcripts by adapting a pre-trained text representation model using CNNs, and a non-semantic encoder, which aligns speech representations with low-level acoustic features from the PASE+ model [23]. The outputs of both encoders are combined, and a lightweight classification head is then trained to perform emotion recognition across multiple downstream datasets. The key contributions of our work are as follows:-
  • Proposing a novel self-supervised model for speech emotion recognition (SER) consisting of dual encoders: a semantic encoder and an acoustic encoder. Both encoders, being trained using distillation-based losses, enable parameter efficient learning.
  • Developing an adaptation strategy for aligning pre-trained text models with speech inputs using convolutional adapters.
  • Experimenting on eight benchmark speech datasets with diverse SER tasks, showcasing the effectiveness of CARE.
  • Identifying the individual and collective impact of semantic and acoustic representations for emotion recognition.
@@ II-A Audio Feature Extraction for SER
Recently, deep learning-based representations have gained popularity as low-level acoustic features. Notable examples include the SincNet architecture by Ravanelli et al.[24] and interpretable Gaussian filters by Agrawal et al.[25]. The LEAF front-end [26], was utilized by Dutta et al. [11] for speech emotion classification. Typically, these models require end-to-end training of both feature extractors and classifiers to generate meaningful speech representations for downstream tasks. In contrast, the proposed CARE architecture is a self-supervised model designed to generalize across diverse datasets without needing to fine-tune the pre-trained weights.
II-B Self-supervision for SER
One of the earliest self-supervised model for the task of speech emotion recognition was proposed by Pascual et al. [27]. This consisted of processing a speech signal by the SincNet model [24] followed by trainable convolutional blocks to predict a number of speech features such as the waveform, mel-frequency cepstral coefficients (MFCCs), pitch etc. Ravanelli et al. [23] further modified this model by adding more self-supervised tasks such as predicting FBANK and Gammatone features [28] to develop the PASE+ model.
Among the general purpose speech SSL models that were proposed over the years, WavLM [16], was shown to outperform other models such as HuBERT [15] and wav2vec2.0 [14] for emotion recognition. Vesper [20] used a modified masking strategy to emphasize high pitch/energy regions of speech—known indicators of emotion—and derived targets for these masked regions from a WavLM teacher model. This approach allowed to learn a smaller student model with enhanced performance for SER. A similar strategy was employed by Ma et al. in emotion2vec [21], which utilized a pre-trained data2vec model as the teacher. Emotion2vec [21] also learns a global embedding to enhance SER performance. In contrast, the proposed CARE model integrates semantic content along with acoustic features to perform emotion recognition. A distillation loss from a text model enables the semantic encoding in a parameter efficient way.
@@ II-C Speech-text Aligned Representations
The alignment of speech and text modalities has received renewed attention for speech representation learning. The SONAR model [29] aligns a speech encoder with textual representations at the utterance level. With the increasing prominence of large language models (LLMs), recent approaches have integrated speech encoders with LLMs. Notably, the SALMONN model by Tang et al.[30] introduced an audio encoder consisting of Whisper model and a music encoder along with the LLaMA language model [31]. Hu et al.[32] proposed WavLLM, combining Whisper and WavLM encoders with the LLaMA model. These LLM-based approaches harness aligned speech-text representations and enable prompt-based applications. However, their substantial model sizes (e.g., $7$B parameters for SALMONN [30]) present significant computational demands for both training and inference. In contrast, the CARE model achieves superior performance on various downstream datasets with a much smaller size of $160$M parameters.
Summary: The landscape of various SER methods is summarized in Fig. 1. We highlight a clear gap in current modeling frameworks: models either prioritize efficiency with limited performance (those in the lower end of the x-axis), or focus on maximizing performance with increased memory and compute requirements (typically based on LLMs). To address this gap, we propose CARE, that combines the computational efficiency of smaller models with the high performance of large-scale systems, thereby providing a superior trade-off between efficiency and performance.
Figure 1: Scatter plot of inference model size (parameters in millions) versus the SER performance (average weighted F1-score over $8$ datasets). CARE is seen to achieve a better trade-off compared to the existing solutions. For more details, refer Tables II and  IV and the associated discussions.
Figure 2: Block diagram of the proposed CARE model. An ASR system is used to generate the transcripts for the pre-training data which is passed through a pre-trained RoBERTa model to generate targets for the semantic encoder. The acoustic encoder of the model is trained with PASE+ features as targets. Blocks in blue indicate either frozen components or those with no learnable parameters. For the semantic encoder the transformer layers are frozen while the convolutional adapters are trained. As the dimension of the output from the acoustic encoder is $768$, a FC layer is attached to match the PASE+ feature dimension of $256$. This FC layer and the average pool block after the semantic encoder are not used during inference.
@@ III-A Background
III-A1 RoBERTa
One of the significant contributions in creating a text representation model was proposed by Devlin et al. [17]. This architecture, called bidirectional encoder representations from transformer (BERT), was trained with two objectives on a corpus of textual data, namely, predicting words masked out in a sentence and to predict whether two sentences semantically follow each other (referred to as next sentence prediction). Liu et. al [33] trained this architecture on a larger corpus of textual data without the next sentence prediction task. This pre-trained model, known as robust optimized BERT approach (RoBERTa), was shown to outperform BERT in a number of downstream tasks.
@@ III-B CARE Model
Having discussed the need for both semantic and acoustic supervision for our model in Sec. I, we propose a dual encoding scheme for modeling speech. These dual encoders — the semantic and acoustic encoders — process the speech signal through distinct supervisory signals suited to their respective objectives. The chosen supervision for each encoder is detailed as follows:
Semantic supervision: An intuitive target for training a model to understand the semantics (or content) of speech would be the corresponding text transcripts. However, obtaining annotated transcripts for large emotional speech pre-training datasets is resource-intensive. Consequently, we assume the pre-training dataset lacks ground-truth text transcripts (even though we have access to it in our case) to develop a methodology that remains applicable even to datasets without human-generated transcriptions. In such a scenario, pre-trained automatic speech recognition (ASR) systems offer an alternative for generating these transcripts, and we use the open-source ASR model Whisper-large-v3 [34] for this task.
Before delving into further details, we wish to discuss a small caveat here. ASR systems have been shown in various studies to exhibit higher word error rates (WER) on emotional speech compared to neutral speech datasets [22]. However, Li et al. [22] show that since the datasets considered in their study consist of many short utterances, the ASR model does not have enough context to generate the correct transcript. This leads us to choose a dataset with podcast recordings as our pre-training corpus. Podcast recordings provide sufficiently long utterances that improve transcript accuracy and offer a broad content variety suitable for pre-training the semantic encoder. Specifically for our pre-training dataset, we observe a WER of $12.53\%$ which is not substantially high.
Given these generated transcripts, the next question is how best to train the semantic encoder. Since the semantic encoder’s purpose is to align the speech signal with its content to facilitate emotion recognition, an ASR-style alignment loss could be applied. However, a sentence level representation for text is more appropriate for the task of emotion recognition as established by Fan et al. [35]. Therefore, we extract contextual word-level embeddings from the transcripts using a pre-trained RoBERTa model[33] and mean-pool these embeddings to obtain a single feature vector representing the entire transcript. These utterance-level embeddings serve as the supervisory signal, or “teacher,” for the semantic encoder in our CARE model. We denote this utterance level embeddings by $\boldsymbol{y}_{text}$.
Acoustic Supervision: While utterance level embeddings serve as the targets for the semantic encoder, frame level descriptors serve as targets for the acoustic encoder. An attentive reader might enquire as to why such a choice is made. Note that mean-pooled speech representations can capture a wide range of properties—such as speaker identity, accent, and language—beyond emotion alone [36]. Moreover, emotion in speech is often contained in the change of different parameters such as pitch, rhythm or energy from one frame to another [6]. Thus a frame-level target is chosen for the acoustic encoder.
A straightforward approach for these frame level acoustic targets would involve masking parts of the speech signal and reconstructing them, an approach adopted in both Vesper [20] and emotion2vec [21], the studies most closely related to ours. However, random masking is less effective for emotion recognition than selectively masking high-energy or high-pitch regions, as demonstrated by Chen et al. [20]. This highlights the role of low-level speech descriptors—such as filter-bank energies and pitch—as they are rich with emotional cues.
Based on these insights, we choose to predict PASE+ features, which encompass filter-bank energies, pitch, and other low-level descriptors essential for capturing emotion. Specifically, we use frame-level PASE+ features with $256$ dimensions as targets for the acoustic encoder in our CARE model. These features are down-sampled by a factor of $2$, producing target descriptors at a frequency of $50$ Hz. We denote the acoustic targets from the PASE+ model by $\boldsymbol{y}_{pase}$.
@@ III-B1 Model Architecture
The speech signal is first processed through a series of convolutional layers designed to produce frame representations every $20$ ms. These are followed by a stack of six transformer layers, forming the common encoder that serves both the acoustic and semantic encoder pathways in our model.
Recall that the target for the semantic encoder is obtained from the mean-pooled pre-trained RoBERTa representations for the generated text transcripts. The semantic encoder is thus expected to align the speech representations with its corresponding generated transcript. This encoder consists of six transformer layers which are initialized with the weights from a pre-trained text representation model. Being trained with textual data, the transformer layers in the semantic encoder do not generalize to speech representations. To address this, we propose a novel adaptation strategy by introducing two 1D-convolutional blocks—one placed before and one after each transformer layer.
The first block adjusts the speech representations from the common encoder to align them with the internal representations expected by the text-based model. The second block refines these representations post-transformer processing. Additionally, following established practice for processing speech in text models [37, 38, 30], the time-resolution of the speech sequence is reduced before processing by the transformer layers in the semantic encoder. Specifically, the convolutional block preceding each transformer layer down-samples the sequence length by a factor of three, while the block following it up-samples it by the same factor. Each convolutional block consists of a single convolutional layer, with a kernel size of $5$, and input and output channels set to $768$ in order to match the dimension of the pre-trained transformer layers. While adaptation of speech SSL models with convolution layers has been explored in prior works [39, 40], adapting pre-trained text models for speech tasks, using convolutional adapters, is explored for the first time in this work. Importantly, the transformer layers themselves are not updated during training. Finally, the semantic encoder’s output representations are average-pooled to produce an utterance-level representation.
The acoustic encoder also consists of six transformer layers, with its output subsequently mapped to $256$ dimensions, using a fully-connected layer, to match the PASE+ feature targets. Figure 2 provides a block diagram of the CARE model.